OpenStudy (anonymous):

Carla remembers that for every pair of pants she purchased, she purchased three t-shirts. She cannot remember how many of each she purchased, but she recalls she purchased 32 items (t-shirts and pants) all together. a. Write an equation using one variable to calculate the number of pants Carla purchased. b. How many t-shirts did Carla purchase? c. Using the prices from Marie’s Boutique, write an expression that can be used to find the total cost of Carla’s purchase before tax. d. What is the total cost of Carla’s purchase if tax is 7.25%?

OpenStudy (anonymous):

The total amount of items she buys is (x + 3x) = 4x where x is the amount of pants. If she buys 1 pair of pants, she'll get 4*1 = 4 items. 1 pair of pants and 3 shirts. So if she buys 32 shirts we know that 4 * x = 32 32/4 = x = 8 pair of pants

OpenStudy (anonymous):

Pants = totalt / 4

OpenStudy (anonymous):

B) T-shirts = Total - Pants = Total - Totalt / 4

OpenStudy (anonymous):

C) We don't have the prices ?

OpenStudy (anonymous):

C) the prices for Marie's Boutique are as follows: dresses: $84.50 each t-shirts $12.50 each handbags $34.99 each pants $42.99 each

OpenStudy (anonymous):

Okey lets say she buys X items. Then the amount of pants, P, will be X / 4. And the amount of t-shirts, T, will be P * 3 (three times as many t-shirts. Lets say x = 32. P = 32 / 4 = 8 T = 8 * 3 = 24 The cost will be: P * $42.99 + T * $12.50 We change T to P * 3 P * $42.99 + (P * 3) * $12.50 For 32 items, P = 8 8 * $42.99 + (8 * 3) * $12.50 = $643.92

OpenStudy (anonymous):

D) Totalt price - the tax. $643.92 - $643.92 * 0.0725 = $597.24 Or the faster way. 100% - 7.25% = 92.75% = 0.9275 Price after tax = $643.92 * 0.9275 = $597.24